# DGS-Net: Distillation-Guided Gradient Surgery for CLIP Fine-Tuning in AI-Generated Image Detection

### v1

@@ 4.2 Orthogonal Suppression
To effectively suppress task-irrelevant components, it is crucial to first characterize these irrelevant features. Since CLIP is trained for text–image alignment, the visual representations produced by its image encoder inherently contain rich semantic information. Prior studies have explored different ways to handle such semantics: some works [yan2024orthogonal, tan2025c2p] exploit them to enhance model generalization, whereas others [yan2025ns, zhang2025towards] seek to eliminate them in order to emphasize low-level artifact cues. To further investigate the influence of semantic features on AI-generated image detection, we convert each image into a textual description and perform binary classification using only the text features extracted by the text encoder. As illustrated in Figure 3, this approach achieves an accuracy of approximately 60%, suggesting that semantic information indeed carries label-related signals. However, the majority of these semantics act as distractors, reflecting dataset-specific shortcuts that hinder generalization. Therefore, we infer that the semantic priors embedded in CLIP’s visual encoder contain substantial harmful knowledge from pre-training, which interferes with reliable detection of AI-generated images.
Therefore, we extract the harmful gradients from the text branch to suppress pre-trained components that are irrelevant to the detection task. As discussed in Sec.˜3, if the text-branch gradient $g_{\text{text}}$ is positive in a given dimension, a small positive perturbation along that coordinate will increase the loss. Consequently, gradient descent will update the corresponding semantic coordinate in the negative direction to reduce the loss. However, this behavior—i.e., relying on semantic coordinates to minimize the loss—often reflects dataset-specific correlations rather than genuine forensic cues, thereby degrading cross-generator generalization. To address this issue, we regard the positive half-space of the text gradient as a harmful direction that should be suppressed, formulated as:
@@
After applying orthogonal suppression, the gradient update direction shifts from $g_{\text{task}}$ to $\tilde{g}$, effectively removing harmful semantic components from the image-task gradient. More importantly, different from existing methods [zhang2025towards, yan2025ns] that directly discard all semantic information, our approach only modifies the gradient update direction to be orthogonal to the harmful semantic component. In this way, we retain beneficial semantic cues while simultaneously improving the model’s generalization ability.
Figure 3: Classification using only text descriptions achieves an accuracy of approximately 60%. We used BLIP to convert images into textual descriptions and trained the detector directly on these text inputs. The resulting detection accuracy fluctuated around 60%, indicating that semantic information indeed carries label-related signals. However, most of them act as distractors, reflecting dataset-specific shortcuts that hinder cross-generator generalization.
Detection Method
Real
Image
Generative Adversarial Networks Other Diffusion Models mAcc.
Pro-
GAN
Cycle-
GAN
Big-
GAN
Style-
GAN
Style-
GAN2
Gau-
GAN
Star-
GAN
WFIR
Deep-
fake
SDv1.4
SDv1.5
ADM
GLIDE
Mid-
journey
Wukong
VQDM
DALLE2
CNN-Spot [wang2020cnn] 99.0 95.3 18.7 1.8 36.5 22.0 2.5 23.1 1.1 29.3 55.9 55.6 1.8 4.8 5.2 27.6 0.7 4.5 29.0
UnivFD [ojha2023towards] 92.3 98.9 90.5 79.2 55.7 48.7 91.1 96.9 92.8 26.9 96.3 96.0 12.7 75.6 61.2 84.7 45.6 62.3 72.7
FreqNet [tan2024frequency] 89.9 99.4 57.1 51.0 75.1 67.5 9.9 88.4 95.4 35.8 99.9 99.8 37.7 78.9 80.8 98.0 34.1 88.8 71.7
NPR [tan2024rethinking] 99.3 98.9 29.3 16.5 67.1 58.7 1.7 6.5 7.9 0.1 100.0 99.9 26.5 69.2 71.0 97.7 15.4 89.8 53.1
Ladeda [cavia2024real] 99.8 99.7 7.2 29.0 95.6 98.3 4.9 0.0 19.2 0.1 100.0 99.9 27.3 79.7 88.8 97.9 15.5 92.4 58.6
AIDE [yan2024sanity] 93.0 95.3 88.0 96.9 89.6 97.0 90.0 97.0 42.9 9.2 99.8 99.7 92.4 98.7 63.7 98.9 90.1 98.9 85.6
C2P-CLIP* [tan2025c2p] 99.8 100.0 99.8 99.8 72.2 78.0 87.7 100.0 34.9 67.6 100.0 99.9 50.5 73.3 28.4 99.7 93.3 98.1 82.4
DFFreq [yan2025generalizable] 98.0 98.0 70.9 93.6 99.0 98.9 93.8 97.4 3.5 76.0 99.9 99.8 65.9 87.8 94.8 99.1 76.0 95.9 85.2
SAFE [li2025improving] 99.2 99.7 85.6 83.5 88.1 96.7 89.0 99.9 8.4 17.3 99.8 99.6 36.8 90.5 86.3 98.5 84.0 92.0 80.8
Effort [yan2024orthogonal] 99.8 100.0 100.0 100.0 79.6 82.1 100.0 100.0 99.9 76.6 100.0 99.9 53.1 81.2 44.2 99.9 96.3 75.8 88.1
NS-Net [yan2025ns] 97.7 100.0 99.9 100.0 95.7 98.6 98.5 100.0 68.9 60.6 100.0 99.9 85.3 97.2 77.4 100.0 98.9 98.8 93.2
Table 1: Cross-model Accuracy (Acc.) Performance on the GenImage [zhu2024genimage] Dataset. The first column represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
Pro-
GAN
Cycle-
GAN
Big-
GAN
Style-
GAN
Style-
GAN2
Gau-
GAN
Star-
GAN
WFIR
Deep-
fake
SDv1.4
SDv1.5
ADM
GLIDE
Mid-
journey
Wukong
VQDM
DALLE2
CNN-Spot[wang2020cnn] 99.9 91.0 59.2 94.8 94.2 86.1 82.5 65.5 74.2 97.6 97.6 66.6 83.1 80.4 88.9 57.0 87.8 84.3
UnivFD [ojha2023towards] 99.9 96.2 94.7 90.9 91.1 97.6 99.7 83.0 83.5 99.3 99.1 65.2 95.3 91.7 97.4 87.2 89.8 91.9
FreqNet [tan2024frequency] 100.0 97.5 87.9 95.9 94.2 54.0 100.0 57.0 88.5 99.9 99.9 73.9 94.1 94.4 99.1 74.9 92.3 88.4
NPR [tan2024rethinking] 100.0 98.6 79.4 98.6 99.6 71.3 97.6 65.5 92.5 100.0 99.9 74.1 97.4 97.8 99.9 81.5 99.3 91.4
Ladeda [cavia2024real] 100.0 99.0 89.5 100.0 100.0 96.4 90.3 93.6 92.6 95.5 95.8 84.8 96.6 93.5 93.0 84.5 93.9 94.0
AIDE [yan2024sanity] 99.6 98.2 91.6 99.4 99.9 74.1 99.7 90.8 66.5 99.9 99.9 99.4 99.9 90.3 99.9 98.9 99.9 94.5
C2P-CLIP* [tan2025c2p] 100.0 100.0 99.9 99.3 99.7 98.8 100.0 99.6 98.3 100.0 100.0 96.3 98.9 90.4 100.0 99.9 100.0 98.8
DFFreq [yan2025generalizable] 100.0 98.7 98.4 100.0 100.0 98.1 99.8 89.0 87.5 100.0 99.9 98.7 99.5 99.7 100.0 99.3 99.9 98.1
SAFE [li2025improving] 100.0 99.5 97.8 99.9 100.0 96.6 100.0 73.8 91.0 100.0 100.0 98.1 99.9 99.9 100.0 99.9 100.0 97.4
Effort [yan2024orthogonal] 100.0 100.0 100.0 99.3 99.2 100.0 100.0 91.1 97.9 100.0 100.0 97.5 99.8 98.9 100.0 100.0 99.9 99.0
NS-Net [yan2025ns] 100.0 100.0 99.7 99.8 99.9 97.0 100.0 99.5 95.2 100.0 100.0 100.0 99.8 97.0 100.0 100.0 100.0 99.2
Table 2: Cross-model Average Precision (A.P.) Performance on the GenImage [zhu2024genimage] Dataset.
@@ 4.3 Prior Alignment
Similar to the approach described in Sec.˜4.2, we define the negative direction of the image gradient as the beneficial direction, expressed as:
@@ 4.4 Overall Process and Loss Function
Real Image 98.2 73.3 64.6 93.8 91.7 88.1 93.8 99.3 92.4 96.9 88.3 83.0
ProGAN 95.3 98.9 99.4 98.9 99.7 95.3 100.0 78.4 99.7 100.0 100.0 100.0
@@
Average 15.0 71.5 66.6 43.9 56.4 67.8 51.5 58.8 64.7 58.7 67.4 81.6
Table 3: Cross-model Accuracy (Acc.) Performance on the AIGIBench [li2025artificial] Dataset. The first row represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
@@ 5.1 Settings
Training datasets. We adopt the Setting-II of AIGIBench [li2025artificial], which includes 144k images generated by ProGAN [karras2018progressive] and SDv1.4 [rombach2022high], along with real images.
Testing datasets. To comprehensively evaluate the effectiveness of our method, we use three benchmarks for detection analysis, including UniversalFakeDetect (shown in Sec.˜8 of the Appendix), GenImage, and AIGIBench. The UniversalFakeDetect [ojha2023towards] contains 8 subsets from various generative models, including Guided [dhariwal2021diffusion], GLIDE [nichol2021glide], LDM [rombach2022high] and DALLE [ramesh2021zero]. The GenImage [zhu2024genimage] datasets comprise 17 generators, including ProGAN [karras2018progressive], CycleGAN [zhu2017unpaired], BigGAN [brock2018large], StyleGAN [karras2019style], StyleGAN2 [karras2020analyzing], GauGAN [park2019semantic], StarGAN [choi2018stargan], WFIR [WFIR], Deepfake [rossler2019faceforensics++], SDv1.4 [rombach2022high], SDv1.5 [rombach2022high], ADM [dhariwal2021diffusion], GLIDE [nichol2021glide], Midjourney [midjourney], Wukong [wukong], VQDM [gu2022vector] and DALLE2 [ramesh2022hierarchical]. The AIGIBench [li2025artificial] includes more comprehensive and advanced generative models, including ProGAN [karras2018progressive], R3GAN [huang2024gan], StyleGAN3 [karras2021alias], StyleGAN-XL [sauer2022stylegan], StyleSwim [zhang2022styleswin], WFIR [WFIR], BlendFace[shiohara2023blendface], E4S [liu2023fine], FaceSwap [Faceswap], InSwap [inswapper], SimSwap [chen2020simswap], DALLE-3 [dalle3], FLUX1-dev [FLUX_1], Midjourney-V6 [midjourney], GLIDE [nichol2021glide], Imagen3 [Imagen_3], SD3 [esser2024scaling], SDXL [podell2023sdxl], BLIP [li2022blip], Infinite-ID [wu2024infinite], PhotoMaker [li2024photomaker], Instant-ID [wang2024instantid], IP-Adapter [ye2023ip], CommunityAI and SocialRF.
Implementation Details. Our method is implemented using the PyTorch framework on an NVIDIA RTX 4090 GPU. The model is trained with the Adam optimizer [kingma2014adam], a learning rate of $1\times 10^{-4}$, and a batch size of 32 for only 1 epoch. The ViT-L/14 model of CLIP is adopted as the pre-trained model [ojha2023towards]. In addition, we add Low-Rank Adaptation to fine-tune the CLIP’s image encoder, following [zanella2024low]. We configure the hyperparameters of the LoRA layers as follows: lora_r = 6, lora_alpha = 6, and lora_dropout = 0.8. The hyperparameter $\lambda$ is set to 0.2. We use Patch Selection [yan2025ns] as data preprocessing to resize the image to $224\times 224$. Importantly, to ensure a fair comparison, we retrain all baseline methods  [^1] When training C2P-CLIP [tan2025c2p], we replace the original text generation module (ClipCap) with BLIP, keeping all other components unchanged. on the same training set described above. More details are shown in Sec.˜7 of the Appendix.
@@ 5.2 Comparison with the State-of-the-Art
Evaluation on GenImage. The results of accuracy (Acc.) and average precision (A.P.) are shown in Table 1 and 2 respectively. Overall, our method achieves 97.6% mAcc. and 99.8% mA.P. across 17 test subsets. Although inference remains identical to C2P-CLIP, our method improves 15.2% mAcc. and 1.0% mA.P., indicating a substantial improvement in CLIP-based detection. Furthermore, compared with the state-of-the-art, NS-Net, our method achieves an additional 4.4% improvement in accuracy. Notably, on the challenging Deepfake dataset, it reaches 96.7% in accuracy, whereas existing methods have historically shown very low performance on this dataset.
Evaluation on AIGIBench. We show the results of accuracy (Acc.) on the AIGIBench dataset in Table 3. The additional metric about A.P. is shown in the Sec.˜9 of the Appendix. Our method achieves 81.6% accuracy, representing a 10.1% improvement over the state-of-the-art, UnivFD. Notably, for Deepfake datasets such as BlendFace, where existing detection methods almost completely fail, our method delivers a substantial gain with an average accuracy improvement of about 50%. Nevertheless, performance remains poor on certain datasets (DALLE-3, SocialRF, CommunityAI), similar to other methods. This may be attributed to unknown post-processing operations, which remain an open challenge.
5.3 Ablation Studies
To further assess the effectiveness of our network, we perform ablation studies on its core components, including Orthogonal Suppression and Prior Alignment. The corresponding results are shown in Table 4. Compared with the baseline, our method achieves 17.7% and 29.2% improvements in accuracy on the GenImage [zhu2024genimage] and AIGIBench [li2025artificial] datasets, respectively. This demonstrates that our method substantially enhances CLIP’s ability to extract more generalizable artifact representations for universal AI-generated image detection. Moreover, the inclusion of any single component yields measurable performance gains, further demonstrating the effectiveness of our approach in addressing the challenges of catastrophic forgetting. More ablation studies are shown in Sec.˜10 of the Appendix.
Orthogonal
Suppression
Prior
Alignment
GenImage [zhu2024genimage] AIGIBench [li2025artificial]
@@ 5.3 Ablation Studies
Table 4: Ablation Study of our method. The accuracy averaged on the GenImage [zhu2024genimage] and AIGIBench [li2025artificial] datasets.
@@
CNN-Spot [wang2020cnn] 59.0 59.9 60.5 62.4 63.2 63.3
UnivFD[ojha2023towards] 76.4 74.2 73.8 80.2 76.8 76.3
NPR [tan2024rethinking] 72.8 70.7 69.2 80.6 79.6 81.0
AIDE [yan2024sanity] 74.6 74.1 70.3 88.6 75.8 81.1
SAFE [li2025improving] 57.7 59.6 60.4 88.5 80.6 83.0
NS-Net [yan2025ns] 85.6 82.3 79.6 88.8 86.1 85.9
@@
Table 5: Robustness on JPEG Compression and Gaussian Blur of Our Method. The accuracy (%) averaged on GenImage [zhu2024genimage].
5.4 Robustness Evaluation
In real-world scenarios, images are inevitably affected by unknown disturbances during transmission and interaction, which pose additional challenges for AI-generated image detection. To investigate robustness under such conditions, we further evaluate the performance of different detection methods against a variety of disturbances, such as JPEG compression and Gaussian blur. As illustrated in Table 5 and 6, experiments on two different datasets show that all methods suffer performance degradation under perturbations, which weakens the detectors’ ability to reliably distinguish real from fake images. Despite these challenging conditions, our method consistently outperforms competing approaches, maintaining relatively high accuracy. Notably, our method exhibits almost no degradation in detection performance under varying levels of Gaussian blur interference, further demonstrating its strong robustness.
@@ 5.4 Robustness Evaluation
CNN-Spot [wang2020cnn] 53.8 54.3 54.5 56.5 56.5 56.6
UnivFD[ojha2023towards] 65.6 66.3 65.1 72.2 68.6 68.6
NPR [tan2024rethinking] 59.0 58.9 58.6 62.9 62.0 62.9
AIDE [yan2024sanity] 61.0 59.6 55.9 75.5 63.4 65.2
SAFE [li2025improving] 51.3 47.9 47.8 75.5 67.6 69.3
NS-Net [yan2025ns] 68.9 67.1 64.9 72.8 70.3 70.2
@@
Table 6: Robustness on JPEG Compression and Gaussian Blur of Our Method. The accuracy (%) averaged on AIGIBench [li2025artificial].

### v2

@@ 4.2 Orthogonal Suppression
To effectively suppress task-irrelevant components, it is crucial to first characterize these irrelevant features. Large-scale pre-trained models exhibit strong semantic representation capabilities, as evidenced by their performance on ImageNet classification. Prior studies have explored different ways to handle such semantics: some works (yan2024orthogonal; tan2025c2p) exploit them to enhance model generalization, whereas others (yan2025ns; zhang2025towards) seek to eliminate them in order to emphasize low-level artifact cues. To further investigate the influence of semantic features on AI-generated image detection, we convert each image into a textual description and perform binary classification using only the text features extracted by the text encoder. As illustrated in Figure 3, this approach achieves an accuracy of approximately 60%, suggesting that semantic information is partially correlated with the labels. However, the majority of these semantics act as distractors, which may hinder generalization. Therefore, we explicitly characterize the semantic priors embedded in CLIP’s visual encoder and only suppress the harmful semantic components from pre-training, which interfere with the reliable detection of AI-generated images.
Specifically, we extract the harmful gradients from the text branch to suppress pre-trained components that are irrelevant to the detection task. As discussed in Section˜3, if the text-branch gradient $g_{\text{text}}$ is positive in a given dimension, a small positive perturbation along that coordinate will increase the loss. Consequently, gradient descent will update the corresponding semantic coordinate in the negative direction to reduce the loss. However, some update directions, which are often associated with detection errors, may act as harmful directions that conflict with genuine forensic cues, ultimately impairing cross-generator generalization. To address this issue, we regard the positive half-space of the text gradient as a harmful direction that should be suppressed, formulated as:
@@
After applying orthogonal suppression, the gradient update direction shifts from $g_{\text{task}}$ to $\tilde{g}$, effectively removing harmful semantic components from the image-task gradient. More importantly, different from existing methods (zhang2025towards; yan2025ns) that directly discard all semantic information, our approach only modifies the gradient update direction to be orthogonal to the harmful semantic component. In this way, we retain beneficial semantic cues while simultaneously improving the model’s generalization.
@@ 4.3 Prior Alignment
Similar to the approach described in Section˜4.2, we define the negative direction of the image gradient as the beneficial direction, expressed as:
@@ 4.4 Overall Process and Loss Function
Table 3: Cross-model Accuracy (Acc.) Performance on the AIGIBench (li2025artificial) Dataset. The first row represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
@@
Real Image 98.2 73.3 64.6 93.8 91.7 88.1 93.8 91.8 92.4 96.9 88.3 83.0
ProGAN 95.3 98.9 99.4 98.9 99.7 95.3 100.0 99.3 99.7 100.0 100.0 100.0
@@
Average 15.0 71.5 66.6 43.9 56.4 67.8 51.5 59.3 64.7 58.7 67.4 81.6
@@ 5.1 Settings
Training datasets. We adopt the Setting-II of AIGIBench (li2025artificial), which includes 144k images generated by ProGAN (karras2018progressive) and SDv1.4 (rombach2022high), along with real images.
Testing datasets. To comprehensively evaluate the effectiveness of our method, we use three benchmarks called UniversalFakeDetect (ojha2023towards) (Appendix˜B of the Appendix), AIGCDetectBench (zhong2023patchcraft), and AIGIBench (li2025artificial), containing Guided (dhariwal2021diffusion), GLIDE (nichol2021glide), LDM (rombach2022high), DALLE (ramesh2021zero), ProGAN (karras2018progressive), CycleGAN (zhu2017unpaired), BigGAN (brock2018large), StyleGAN (karras2019style), StyleGAN2 (karras2020analyzing), GauGAN (park2019semantic), StarGAN (choi2018stargan), WFIR (WFIR), Deepfake (rossler2019faceforensics++), SDv1.4 (rombach2022high), SDv1.5 (rombach2022high), ADM (dhariwal2021diffusion), Midjourney (midjourney), Wukong (wukong), VQDM (gu2022vector), DALLE2 (ramesh2022hierarchical), R3GAN (huang2024gan), StyleGAN3 (karras2021alias), StyleGAN-XL (sauer2022stylegan), StyleSwim (zhang2022styleswin), BlendFace(shiohara2023blendface), E4S (liu2023fine), FaceSwap (Faceswap), InSwap (inswapper), SimSwap (chen2020simswap), DALLE-3 (dalle3), FLUX1-dev (FLUX_1), Imagen3 (Imagen_3), SD3 (esser2024scaling), SDXL (podell2023sdxl), BLIP (li2022blip), Infinite-ID (wu2024infinite), PhotoMaker (li2024photomaker), Instant-ID (wang2024instantid), IP-Adapter (ye2023ip), CommunityAI and SocialRF.
Implementation Details. The model is trained with the Adam optimizer (kingma2014adam), a learning rate of $1\times 10^{-4}$, and a batch size of 32 for only 1 epoch. The ViT-L/14 model of CLIP is adopted as the pre-trained model (ojha2023towards). In addition, we add Low-Rank Adaptation to fine-tune the CLIP’s image encoder, following (zanella2024low). The hyperparameter $\lambda$ is set to 0.2. We use Patch Selection (yan2025ns) to resize the image to $224\times 224$. Importantly, to ensure a fair comparison, we retrain all baseline methods  [^1] When training C2P-CLIP (tan2025c2p), we replace the original text generation module (ClipCap) with BLIP, keeping all other components unchanged. on the same training set. More details are shown in Appendix˜A of the Appendix.
@@ 5.2 Comparison with the State-of-the-Art
Evaluation on AIGCDetectBench. The results of accuracy (Acc.) and average precision (A.P.) are shown in Table 1 and 2 respectively. Overall, our method achieves 97.6% mAcc. and 99.8% mA.P. across 17 test subsets. Although inference remains identical to C2P-CLIP, our method improves 15.2% mAcc. and 1.0% mA.P., indicating a substantial improvement in CLIP-based detection. Furthermore, compared with the state-of-the-art NS-Net, our method achieves an additional 4.4% improvement in mAcc. Notably, on the challenging Deepfake dataset, it reaches 96.7% in accuracy, whereas existing methods have historically shown very low.
Evaluation on AIGIBench. We show the results of accuracy (Acc.) on the AIGIBench dataset in Table 3. The additional metric about A.P. is shown in the Appendix˜C of the Appendix. Our method achieves 81.6% accuracy, a 10.1% improvement over the state-of-the-art, UnivFD. Notably, for Deepfake datasets such as BlendFace, where existing detection methods almost completely fail, our method delivers a substantial gain with an average accuracy improvement of about 50%. Nevertheless, performance remains poor on certain datasets (DALLE-3, SocialRF, CommunityAI), similar to other methods. This may be attributed to unknown post-processing operations, remaining an open challenge.
Table 4: Ablation Study of our method. The accuracy averaged on the AIGCDetectBench and AIGIBench datasets.
Orthogonal Suppression Prior Alignment AIGCDetectBench AIGIBench
@@
Table 5: Robustness on JPEG Compression and Gaussian Blur of Our Method. The accuracy averaged on AIGCDetectBench.
@@
CNN-Spot (wang2020cnn) 59.0 59.9 60.5 62.4 63.2 63.3
UnivFD(ojha2023towards) 76.4 74.2 73.8 80.2 76.8 76.3
NPR (tan2024rethinking) 72.8 70.7 69.2 80.6 79.6 81.0
AIDE (yan2024sanity) 74.6 74.1 70.3 88.6 75.8 81.1
SAFE (li2025improving) 57.7 59.6 60.4 88.5 80.6 83.0
NS-Net (yan2025ns) 85.6 82.3 79.6 88.8 86.1 85.9
@@
Table 6: Robustness on JPEG Compression and Gaussian Blur of Our Method. The accuracy averaged on AIGIBench.
@@
CNN-Spot (wang2020cnn) 53.8 54.3 54.5 56.5 56.5 56.6
UnivFD(ojha2023towards) 65.6 66.3 65.1 72.2 68.6 68.6
NPR (tan2024rethinking) 59.0 58.9 58.6 62.9 62.0 62.9
AIDE (yan2024sanity) 61.0 59.6 55.9 75.5 63.4 65.2
SAFE (li2025improving) 51.3 47.9 47.8 75.5 67.6 69.3
NS-Net (yan2025ns) 68.9 67.1 64.9 72.8 70.3 70.2
@@
5.3 Ablation Studies
To further assess the effectiveness of our network, we perform ablation studies on its core components. The corresponding results are shown in Table 4. Compared with the baseline, our method achieves 17.7% and 29.2% improvements in accuracy on the AIGCDetectBench and AIGIBench datasets, respectively. This demonstrates that our method substantially enhances CLIP’s ability to extract more generalizable artifact representations for universal AI-generated image detection. Moreover, the inclusion of any single component yields measurable performance gains, further demonstrating the effectiveness of our approach in addressing the challenges of catastrophic forgetting. More ablation studies are shown in Appendix˜D of the Appendix.
5.4 Robustness Evaluation
In real-world scenarios, images are inevitably affected by unknown disturbances during transmission and interaction. To investigate robustness under such conditions, we further evaluate the performance of different detection methods against a variety of disturbances, such as JPEG compression and Gaussian blur. As illustrated in Table 5 and 6, our method consistently outperforms competing approaches, maintaining relatively high accuracy. Notably, our method exhibits almost no degradation in detection performance under varying levels of Gaussian blur interference, further demonstrating its strong robustness.